OpenStudy (anonymous):

You have 7/8 cup carrots,1/3 cup onions,and 1/2 cup red peppers. About how many cups of vegetables do you have in all? a) about 1 cup b) about 2 cups c) about 3 cups d) about 1/2 cup

OpenStudy (anonymous):

@GhostBirdSeriesNerd

OpenStudy (anonymous):

I will give medal

OpenStudy (anonymous):

so first, u'll have to convert the fractions so that all of the denominators are the same

OpenStudy (anonymous):

lets make the denominator 48, since 8*3*2 equals 48

OpenStudy (anonymous):

7/8+1/2+1/3=

OpenStudy (anonymous):

the first fraction would thereby be 42/48, the second would be 16/48, and the last would be 24/48

OpenStudy (anonymous):

you need common denominator

OpenStudy (anonymous):

so the common should be 24

OpenStudy (anonymous):

now just add the numerators, keep the same denominator, divide, round, and bingo! :)

OpenStudy (anonymous):

ps the answer is 1.7083, so now just round

OpenStudy (anonymous):

7/8*3/3=21/24 1/2*12/12=12/24 1/3*8/8=8/24

OpenStudy (anonymous):

don't listen to them it is me tbing

OpenStudy (anonymous):

What!!!!

OpenStudy (anonymous):

1.8000

OpenStudy (anonymous):

well either way the ending answer is the same, both round to 2

OpenStudy (anonymous):

(21+12+8)/24=41/24

OpenStudy (anonymous):

Thank you @desiswag , @AJ01

OpenStudy (anonymous):

np :)